### v1

# on his 75th birthday with deep respect and admiration

Abstract
We discuss the Dirac oscillator in $(1+1)$ and $(2+1)$ dimensions and present its generalizations in the spirit of the isotonic oscillator. In $(1+1)$ dimensions, the Dirac oscillator returns to quantum harmonic oscillator in the non-relativistic limit, while its generalization maps to the isotonic oscillator. We describe exact solutions of these generalized systems and also present their non-relativistic limits. Finally, based on supersymmetric quantum mechanics, we show that a generalized Dirac oscillator in $(2+1)$ dimensions can be mapped to an anti-Jaynes-Cummings-like Hamiltonian in which the spin operators couple with the supercharges.
I Introduction
The Dirac equation (see for example, dirac ) is one of the most important discoveries of the 20th century. It has been solved exactly in multiple cases, including the free particle and the hydrogen atom. Further, it has appeared frequently in recent literature, mainly in the context of graphene as the (electronic) dispersion relation of the latter can take a form linear in the momentum graphene . Itô and coworkers diracosc1 studied the Dirac equation by adding a linear term $-imc\omega\beta{{\mbox{\boldmath$\alpha$}}}{\bf\cdot r}$ and subsequently it was observed that in the non-relativistic limit, the system reduces to a harmonic oscillator with a strong spin-orbit coupling and so referred to as the Dirac oscillator by Moshinsky and Szczepaniak diracosc (see also, diracoscquesne ). The Dirac oscillator has been studied extensively from different viewpoints including supersymmetric quantum mechanics susyquesne ; susydo (see also, susydirac ; bagchidirac ), magnetism diracB ; app3 , and quantum phase transitions qpt ; qpt1 . Moreover, the Dirac oscillator has attracted a lot of attention and has found many physical applications in various areas of physics app7 ; app6 ; app10 ; app1 ; app4 ; app8 ; app9 ; app2 ; app5 ; ext1 ; ext2 . An exciting property of the Dirac oscillator lies in its connection with quantum optics opt ; opt1 ; opt2 , where in $(2+1)$ dimensions it maps to the Jaynes-Cummings/anti-Jaynes-Cummings models qpt1 ; diracB , describing the atomic transitions in a two-level system.
The harmonic oscillator plays a key role in classical and quantum mechanics, mainly due to the fact that it is an exactly-solvable system but also due to its ubiquitous appearance in different problems of physical interest, from mechanics to quantum field theory. An intriguing aspect of the classical harmonic oscillator is the insensitivity of the period of oscillation to the amplitude or energy, i.e., the system admits a constant frequency, irrespective of the maximum displacement from the origin. Such systems are called ‘isochronous’ and have been studied in detail by several authors urabe ; CV ; Calo ; PG_AGC_symp ; PG_AGC_iso ; particularly notable amongst these is the work of Urabe urabe . With isochronous systems being rather rare, it has been demonstrated that for one-dimensional systems admitting potentials that are rational functions of the position variable, there are only two potentials that support isochronous oscillations CV :
$V_{\rm ho}(x)=\frac{m\omega^{2}x^{2}}{2},\quad\quad V_{\rm io}(x)=Ax^{2}+\frac{B}{x^{2}},$ (1)
where $m,\omega,A,B>0$ are constants. While $V_{\rm ho}(x)$ which is the harmonic potential is defined on the entire real line, the other one, i.e., $V_{\rm io}(x)$, dubbed the isotonic potential is defined on the half-line, i.e., either for $x>0$ or for $x<0$ due to the quadratic singularity at $x=0$; for definiteness, we will take $x>0$. Various other isochronous dynamical systems can often be mapped to either the harmonic oscillator or the isotonic oscillator (see for example, Lie1 ; Lie2 ; mom ). It should be mentioned that the work of Calogero and coworkers (see for example, Calo ) has demonstrated that isochronous systems are not as rare as their existence was initially thought to be (see also, PG_AGC_symp ). Therefore, apart from the harmonic oscillator whose importance is well known, the isotonic oscillator presents yet another intriguing exactly-solvable problem with properties similar to the former. The classical (Newtonian) equation of motion arising from such a potential is well known as the Ermakov-Pinney equation EP and arises in the context of the Ermakov-Lewis invariant of the time-dependent harmonic oscillator (see paddy and references therein).
Quantum mechanically, isochronous systems are associated with an equispaced spectra (in non-relativistic quantum mechanics) iso_car . In fact, the isotonic oscillator is exactly solvable and admits a spectrum of equispaced energy levels iso1 ; iso2 . Due to the importance of the isotonic potential as asserted above, it seems worthwhile to explore the extension of the Dirac oscillator in the spirit of the isotonic potential. Although some preliminary results for the same in $(1+1)$ dimensions appear in bagchidirac , in this paper, we propose exactly-solvable extensions of the Dirac equation in the spirit of the isotonic oscillator in both $(1+1)$ and $(2+1)$ dimensions and also treat their harmonic counterparts. Interestingly, for the $(2+1)$-dimensional extension, the non-relativistic limit admits an intriguing angular-momentum term. Finally, based on supersymmetric quantum mechanics and the known connection between the Dirac oscillator and the Hamiltonian of a two-level system interacting with a quantized single-mode (optical) field, we will propose a generalization of the same in which the two-level system couples with the supercharge operators. With this brief introduction, let us begin our formal analysis with the $(1+1)$-dimensional case; the $(2+1)$-dimensional case will be taken up subsequently.
II (1+1)-dimensional case
Let us begin with the $(1+1)$-dimensional Dirac Hamiltonian
$H=c\alpha(p_{x}-i\beta W(x))+\beta mc^{2},$ (2)
$\alpha$ and $\beta$ are 2 $\times$ 2 matrices, $W(x)$ is a real-valued function that we will specify below, and other symbols have their usual meanings. For our purposes, we have $\alpha=\sigma_{x}$ and $\beta=\sigma_{z}$, where $\sigma_{x}$, $\sigma_{y}$, and $\sigma_{z}$ are the Pauli matrices. Thus, in the matrix form, the Hamiltonian turns out to be
$H=\begin{pmatrix}mc^{2}&c(p_{x}+iW(x))\\ c(p_{x}-iW(x))&-mc^{2}\end{pmatrix},$ (3)
which is obviously Hermitian for a real-valued function $W(x)$. Taking
$\Psi=\begin{pmatrix}\psi_{1}\\ \psi_{2}\end{pmatrix}$ (4)
to be a time-independent eigenfunction of $H$ (the Dirac wavefunction), using $H\Psi=E\Psi$ for real $E$, one gets the following first-order equations:
$c(p_{x}-iW(x))\psi_{1}$ $=$ $(E+mc^{2})\psi_{2},$ (5)
$c(p_{x}+iW(x))\psi_{2}$ $=$ $(E-mc^{2})\psi_{1}.$ (6)
Considering (6), multiplying both sides with $(E+mc^{2})$ and upon using (5), we get (we will set $\hbar=1$)
$c^{2}\bigg{[}-\frac{d^{2}}{dx^{2}}+W(x)^{2}-W^{\prime}(x)\bigg{]}\psi_{1}=[E^{2}-(mc^{2})^{2}]\psi_{1},$ (7)
where we have used $p=-i(d/dx)$. Writing $E=mc^{2}+\epsilon$, in the non-relativistic limit, i.e., for $c\rightarrow\infty$, we find $E^{2}-(mc^{2})^{2}\approx 2mc^{2}\epsilon$, and this means the equation (7) turns out to be
$\bigg{[}-\frac{1}{2m}\frac{d^{2}}{dx^{2}}+\frac{W(x)^{2}}{2m}-\frac{W^{\prime}(x)}{2m}\bigg{]}\psi_{1}=\epsilon\psi_{1}.$ (8)
One can similarly derive the evolution equation for $\psi_{2}$ which makes the supersymmetric connection clear if we identify $W(x)$ as the superpotential susydirac ; bagchidirac . However, throughout this paper we will explicitly consider only $\psi_{1}$ for the sake of simplicity as $\psi_{2}$ can then be obtained from equation (5) once $\psi_{1}$ is known. In the following, we will refer to $\psi_{1}$ as the wavefunction.
II.1 (1+1)-dimensional Dirac oscillator
In order to observe the $(1+1)$-dimensional Dirac oscillator, the counterpart of the case studied in diracosc (see also, diracB ), let us take
$W(x)=m\omega x,\quad\quad x\in\mathbb{R},$ (9)
where $m,\omega>0$ are constants. Thus, (7) turns out to be
$\bigg{[}-c^{2}\frac{d^{2}}{dx^{2}}+m^{2}c^{2}\omega^{2}x^{2}-mc^{2}\omega\bigg{]}\psi_{1}=[E^{2}-(mc^{2})^{2}]\psi_{1}.$ (10)
The above equation possess the same operator structure as the time-independent Schrödinger equation of a harmonic oscillator. In particular, if we write
$m\omega\bigg{[}-\frac{d^{2}}{du^{2}}+u^{2}\bigg{]}\psi_{1}=\frac{[E^{2}-(mc^{2})^{2}+mc^{2}\omega]}{c^{2}}\psi_{1},$ (11)
where we have defined $u=\sqrt{m\omega}x$, then the solution of the above-mentioned equation gives
$\psi_{1,n}(x)=\frac{\pi^{1/4}}{\sqrt{2^{n}n!}}e^{-m\omega x^{2}/2}H_{n}(\sqrt{m\omega}x),\quad\quad E_{n}^{2}=(mc^{2})^{2}\bigg{[}1+\frac{2n\omega}{mc^{2}}\bigg{]},$ (12)
where $\{H_{n}(\cdot)\}$ are the Hermite polynomials. This is the $(1+1)$-dimensional analogue of the Dirac oscillator described by Moshinsky and Szczepaniak diracosc . In the non-relativistic limit, (8) becomes
$\bigg{[}-\frac{1}{2m}\frac{d^{2}}{dx^{2}}+\frac{m\omega^{2}x^{2}}{2}-\frac{\omega}{2}\bigg{]}\psi_{1}=\epsilon\psi_{1},$ (13)
which is just the harmonic oscillator and can be solved to give the spectrum $\epsilon_{n}=n\omega$. Note that unlike Moshinsky and Szczepaniak diracosc , the non-relativistic limit does not contain a spin-orbit term – this is due to the fact that we have considered the $(1+1)$-dimensional Dirac oscillator unlike the $(3+1)$-dimensional case as considered in diracosc .
II.2 (1+1)-dimensional Dirac-isotonic oscillator
In what follows, we will extend this construction to describe the relativistic analogue of the isotonic oscillator iso1 ; iso2 . For that, let us consider the following form of the function $W(x)$:
$W(x)=ax+\frac{b}{x},\quad\quad x>0,$ (14)
where $a,b>0$ are constants. Thus, the problem is restricted to the real half-line for which $x>0$, i.e., the superpotential is singular at $x=0$ (see also, bagchidirac ; susypani ). The limit $b\rightarrow 0$ recovers the harmonic-oscillator case ($a=m\omega$) but only on the real half-line, although the problem can be analytically continued on the full real line. Now, substituting (14) into equation (7), one gets the following Klein-Gordon equation for $\psi_{1}$:
$\bigg{[}-\frac{d^{2}}{dx^{2}}+a^{2}x^{2}+\frac{b(b+1)}{x^{2}}\bigg{]}\psi_{1}=\frac{[E^{2}-(mc^{2})^{2}+ac^{2}(-2b+1)]}{c^{2}}\psi_{1}.$ (15)
We will say that equation (15) represents a ‘Dirac-isotonic’ oscillator, defined on $x>0$. Since the left-hand side has the same operator structure as the Schrödinger equation, we can easily solve the above-mentioned equation following the treatment presented in iso1 . The wavefunctions turn out to be (up to normalization factors)
$\psi_{1,n}(x)\sim x^{\nu}e^{-\frac{ax^{2}}{2}}{{}_{1}}F_{1}\bigg{(}-n,\nu+\frac{1}{2},ax^{2}\bigg{)},\quad\quad\nu=\frac{1}{2}[\sqrt{1+4b(b+1)}+1],$ (16)
while the spectrum is
$E_{n}^{2}=(mc^{2})^{2}\bigg{[}1+\frac{a}{m}\big{(}4n+2b+1+\sqrt{1+4b(b+1)}\big{)}\bigg{]},$ (17)
where $n=0,1,2,\cdots$. The function ${{}_{1}}F_{1}(\cdot,\cdot,\cdot)$ is the confluent hypergeometric function that solves Kummer’s differential equation and because $n$ is a positive integer, it can be expressed in terms of Laguerre polynomials handbook . In the non-relativistic limit, the equation (15) returns to a Schrödinger equation of a particle with mass $m$ in an isotonic potential (with constant shift):
$\bigg{[}-\frac{1}{2m}\frac{d^{2}}{dx^{2}}+\frac{a^{2}x^{2}}{2m}+\frac{b(b+1)}{2mx^{2}}+\frac{a(2b-1)}{2m}\bigg{]}\psi_{1}=\epsilon\psi_{1},$ (18)
where $\epsilon=E-mc^{2}$. In what follows, we will explore the (2+1)-dimensional case.
III (2+1)-dimensional case
Let us now consider the $(2+1)$-dimensional Dirac Hamiltonian
$H=c\alpha_{x}(p_{x}-i\beta W_{x}(\mathbf{r}))+c\alpha_{y}(p_{y}-i\beta W_{y}(\mathbf{r}))+\beta mc^{2},$ (19)
where $\mathbf{r}=(x,y)$, $\alpha_{x}=\sigma_{x}$, $\alpha_{y}=\sigma_{y}$, and $\beta=\sigma_{z}$. The ‘superpotential’ is a vector-valued function of $\mathbf{r}=(x,y)$ with $x$ and $y$ components being given by $W_{x}=W_{x}(\mathbf{r})$ and $W_{y}=W_{y}(\mathbf{r})$, respectively. These are real-valued functions. Explicitly, one can write
$H=\begin{pmatrix}mc^{2}&cP^{*}+ic\mathcal{W}^{*}(\mathbf{r})\\ cP-ic\mathcal{W}(\mathbf{r})&-mc^{2}\end{pmatrix},$ (20)
where $P=p_{x}+ip_{y}$ and $\mathcal{W}(\mathbf{r})=W_{x}(\mathbf{r})+iW_{y}(\mathbf{r})$. Taking the Dirac wavefunction to be of the form dictated by (4) with $\psi_{1,2}=\psi_{1,2}(x,y)$, we get
$c(P^{*}+i\mathcal{W}^{*}(\mathbf{r}))\psi_{2}$ $=$ $(E-mc^{2})\psi_{1},$ (21)
$c(P-i\mathcal{W}(\mathbf{r}))\psi_{1}$ $=$ $(E+mc^{2})\psi_{2}.$ (22)
Focusing on $\psi_{1}$, combining the two first-order equations above, one finds the following second-order equation:
$c^{2}\big{[}p_{x}^{2}+p_{y}^{2}+W_{x}^{2}+W_{y}^{2}+2(W_{y}p_{x}-W_{x}p_{y})-i((p_{x}W_{x})+(p_{y}W_{y}))+((p_{x}W_{y})-(p_{y}W_{x}))\big{]}\psi_{1}=[E^{2}-(mc^{2})^{2}]\psi_{1}.$ (23)
Let us discuss two cases now.
III.1 (2+1)-dimensional Dirac oscillator
Taking $W_{x}=m\omega x$ and $W_{y}=m\omega y$, one finds $(p_{x}W_{y})=0$ and $(p_{y}W_{x})=0$. Thus, one has
$c^{2}\bigg{[}-\bigg{(}\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\bigg{)}+m^{2}\omega^{2}(x^{2}+y^{2})-2m\omega L_{z}-2m\omega\bigg{]}\psi_{1}=[E^{2}-(mc^{2})^{2}]\psi_{1},$ (24)
where $L_{z}=xp_{y}-yp_{x}$. This is the $(2+1)$-dimensional analogue of the Dirac oscillator diracosc . In order to solve it, let us resort to polar coordinates $(r,\theta)$; taking the ansatz $\psi_{1}(r,\theta)=e^{im_{l}\theta}\phi(r)$, we get
$-\frac{d^{2}\phi}{dr^{2}}-\frac{1}{r}\frac{d\phi}{dr}+\bigg{[}m^{2}\omega^{2}r^{2}+\frac{m_{l}^{2}}{r^{2}}\bigg{]}\phi=\frac{[E^{2}-(mc^{2})^{2}+2mc^{2}\omega+2mc^{2}m_{l}\omega]}{c^{2}}\phi,$ (25)
which can be solved using the techniques employed commonly in non-relativistic quantum mechanics. Defining $\Lambda=\frac{[E^{2}-(mc^{2})^{2}+2mc^{2}\omega+2mc^{2}m_{l}\omega]}{c^{2}}$ and $\chi(r)=\sqrt{r}\phi(r)$, one gets the simplified equation
$-\frac{d^{2}\chi}{dr^{2}}+\bigg{[}m^{2}\omega^{2}r^{2}+\frac{m_{l}^{2}-\frac{1}{4}}{r^{2}}\bigg{]}\chi=\Lambda\chi.$ (26)
This looks like (15) and can therefore be solved exactly to give (up to normalization factors)
$\chi_{n}(r)\sim r^{\left(m_{l}+\frac{1}{2}\right)}e^{-\frac{m\omega r^{2}}{2}}{{}_{1}}F_{1}\bigg{(}-n,m_{l}+1,m\omega r^{2}\bigg{)},\quad\quad E_{n}^{2}=(mc^{2})^{2}\bigg{[}1+\frac{4n\omega}{mc^{2}}\bigg{]},$ (27)
where $n=0,1,2,\cdots$. Notice that the full wavefunctions are obtained as $\psi_{1,n}(r,\theta)=r^{-1/2}e^{im_{l}\theta}\chi_{n}(r)$. In the non-relativistic limit ($c\rightarrow\infty$), $E^{2}-(mc^{2})^{2}\approx 2mc^{2}\epsilon$, meaning that one obtains the following Schrödinger equation:
$\bigg{[}-\frac{1}{2m}\bigg{(}\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\bigg{)}+\frac{m\omega^{2}}{2}(x^{2}+y^{2})-\omega L_{z}-\omega\bigg{]}\psi_{1}=\epsilon\psi_{1},$ (28)
and the spectrum simply reads $\epsilon_{n}=\omega\big{(}2n-m_{l}\big{)}$. Thus, in the non-relativistic limit, the problem reduces to the two-dimensional harmonic oscillator but with a ‘spin-orbit’ term that contributes $-\omega m_{l}$ to the spectrum.
III.2 A nonlinear superpotential
Inspired from the choice of superpotential in (14), let us take $W(\mathbf{r})\equiv W_{x}\hat{x}+W_{y}\hat{y}=a\mathbf{r}+\frac{b\mathbf{r}}{|\mathbf{r}|^{2}}$. This implies that
$W_{x}=ax+\frac{bx}{x^{2}+y^{2}},\quad\quad W_{y}=ay+\frac{by}{x^{2}+y^{2}}.$ (29)
Thus, the equation (23) becomes
$\bigg{[}-\bigg{(}\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\bigg{)}+a^{2}(x^{2}+y^{2})+\frac{b^{2}}{x^{2}+y^{2}}-2\bigg{(}a+\frac{b}{x^{2}+y^{2}}\bigg{)}L_{z}\bigg{]}\psi_{1}=\frac{[E^{2}-(mc^{2})^{2}-2ac^{2}(b-1)]}{c^{2}}\psi_{1},$ (30)
which matches with (24) for $a=m\omega$ and $b=0$. It is convenient to employ polar coordinates $(r,\theta)$ on the plane using which, together with the ansatz $\psi_{1}(r,\theta)=e^{im_{l}\theta}\phi(r)$, one finds
$-\bigg{(}\frac{d^{2}\phi}{dr^{2}}+\frac{1}{r}\frac{d\phi}{dr}\bigg{)}+a^{2}r^{2}\phi+\frac{b^{2}+m_{l}(m_{l}-2b)}{r^{2}}\phi=\frac{[E^{2}-(mc^{2})^{2}+2am_{l}c^{2}-2ac^{2}(b-1)]}{c^{2}}\phi.$ (31)
The term with the first derivative can be eliminated by putting $\chi(r)=\sqrt{r}\phi(r)$ and one then finds
$-\frac{d^{2}\chi}{dr^{2}}+a^{2}r^{2}\chi+\frac{(m_{l}-b)^{2}-\frac{1}{4}}{r^{2}}\chi=\Gamma\chi,$ (32)
where we have identified $\Gamma=\frac{[E^{2}-(mc^{2})^{2}+2am_{l}c^{2}-2ac^{2}(b-1)]}{c^{2}}$. The above-mentioned equation has the same operator structure as (15) or (26), and can be solved exactly to give
$\chi_{n}(r)=r^{m_{l}-b+\frac{1}{2}}e^{-\frac{ar^{2}}{2}}{{}_{1}}F_{1}\big{(}-n,m_{l}-b+1,ar^{2}\big{)},\quad\quad E_{n}^{2}=(mc^{2})^{2}\bigg{[}1+\frac{4a}{m^{2}c^{2}}(n+1)\bigg{]},$ (33)
where $n=0,1,2,\cdots$. Notice that the full wavefunctions are obtained as $\psi_{1,n}(r,\theta)=r^{-1/2}e^{im_{l}\theta}\chi_{n}(r)$. Finally, let us remark that the non-relativistic limit is obtained straightforwardly; in particular, in this limit, equation (31) becomes
$\bigg{[}-\frac{1}{2m}\bigg{(}\frac{d^{2}}{dr^{2}}+\frac{1}{r}\frac{d}{dr}\bigg{)}+\frac{a^{2}r^{2}}{2m}+\frac{(m_{l}-b)^{2}}{2mr^{2}}-\frac{am_{l}}{m}+\frac{a(b-1)}{m}\bigg{]}\phi=\epsilon\phi,$ (34)
which resembles the isotonic oscillator for the radial variable. Notice that upon putting $a=m\omega$ and $b=0$, one can exactly recover the (radial) equation representing the non-relativistic limit of the Dirac oscillator.
IV A supersymmetric anti-Jaynes-Cummings model
Since the mapping between the $(2+1)$-dimensional Dirac oscillator and the Jaynes-Cummings/anti-Jaynes-Cummings model opt ; opt1 ; opt2 is well known qpt1 (see also, diracB ), this will allow us to present a generalized anti-Jaynes-Cummings-like Hamiltonian based on the above-studied generalization of the Dirac oscillator. Recall that we arrived at the Dirac-isotonic oscillator via the route based on supersymmetric quantum mechanics. Let us define the supercharge operators
$\mathcal{A}^{\dagger}=\frac{P^{*}+i\mathcal{W}^{*}(\mathbf{r})}{\sqrt{\delta}},\quad\quad\mathcal{A}=\frac{P-i\mathcal{W}(\mathbf{r})}{\sqrt{\delta}},$ (35)
where $P=p_{x}+ip_{y}$, $\mathcal{W}(\mathbf{r})=W_{x}(\mathbf{r})+iW_{y}(\mathbf{r})$, and $\delta$ is a positive constant. It then follows that
$\mathcal{A}^{\dagger}\mathcal{A}\delta=p_{x}^{2}+p_{y}^{2}+W_{x}^{2}+W_{y}^{2}+2(W_{y}p_{x}-W_{x}p_{y})-i((p_{x}W_{x})+(p_{y}W_{y}))+((p_{x}W_{y})-(p_{y}W_{x})),$ (36)
while
$\mathcal{A}\mathcal{A}^{\dagger}\delta=p_{x}^{2}+p_{y}^{2}+W_{x}^{2}+W_{y}^{2}+2(W_{y}p_{x}-W_{x}p_{y})+i((p_{x}W_{x})+(p_{y}W_{y}))+((p_{x}W_{y})-(p_{y}W_{x})).$ (37)
Naturally then, one gets the commutator $[\mathcal{A},\mathcal{A}^{\dagger}]=2i\delta^{-1}((p_{x}W_{x})+(p_{y}W_{y}))$. Taking $W_{x}=m\omega x$ and $W_{y}=m\omega y$ as taken for the $(2+1)$-dimensional Dirac oscillator and upon picking $\delta=4m\omega$, one gets $[\mathcal{A},\mathcal{A}^{\dagger}]=1$, the familiar oscillator algebra. Generally, with nonlinear superpotentials such as that giving the Dirac-isotonic oscillator, one has $[\mathcal{A},\mathcal{A}^{\dagger}]\neq 1$. Nevertheless, employing these operators one can express the Dirac Hamiltonian as
$H=\begin{pmatrix}mc^{2}&c\sqrt{\delta}\mathcal{A}^{\dagger}\\ c\sqrt{\delta}\mathcal{A}&-mc^{2}\end{pmatrix}.$ (38)
Introducing the spin-projection operators $\sigma^{\pm}=\frac{1}{2}(\sigma_{x}\pm\sigma_{y})$, it is simple to show that the Hamiltonian matrix (38) is equivalent to the following form:
$H=g(\sigma^{-}\mathcal{A}+\sigma^{+}\mathcal{A}^{\dagger})+\sigma_{z}mc^{2},$ (39)
where the parameter describing the spin-supercharge coupling is given by $g=c\sqrt{\delta}$. Notice that while the standard anti-Jaynes-Cummings Hamiltonian opt describes the coupling between a two-level system and a single-mode oscillator, the supersymmetric generalization presented above describes coupling between the spin operators (two-level) and the supercharge operators.
V Discussion
Isochronous systems occur in various physical situations and have been widely investigated in the literature. In the present work, we have investigated an isochronous system in the framework of relativistic quantum mechanics by proposing the Dirac-isotonic oscillator in $(1+1)$ and $(2+1)$ dimensions for the first time. This generalized system in $(1+1)$ dimensions reduces to the usual isotonic oscillator in the non-relativistic limit, whereas the Dirac-isotonic oscillator in $(2+1)$ dimensions maps to an isotonic oscillator with strong spin-orbit coupling and an added centrifugal term that couples with the orbital angular momentum. In both $(1+1)$ and $(2+1)$ dimensions, the systems are exactly-solvable ones and in the non-relativistic limit, their spectra are equispaced. We finally described a supersymmetric generalization of the anti-Jaynes-Cummings Hamiltonian starting with defining appropriate supercharge operators for the Dirac Hamiltonian. It would be exciting to explore the possibility of a relationship between the Dirac-isotonic oscillator and optical systems. We keep this issue for future work.
Acknowledgements: We are thankful to Bijan Bagchi for several stimulating discussions. A.G. is grateful to Akash Sinha for discussions and also thanks the Ministry of Education, Government of India for financial support in the form of a Prime Minister’s Research Fellowship (ID: 1200454). A.G. is grateful to the Department of Physics, Banaras Hindu University for hospitality where this work was initiated. B.P.M. acknowledges the incentive research grant for faculty under the IoE Scheme (IoE/Incentive/2021-22/32253) of the Banaras Hindu University.